OpenStudy (anonymous):

Can someone explain why (sqrt3+i)^2 is 4e^i(pi/3) instead of 4e^i(2pi/3) when I try to solve it using De Moivres formula?

OpenStudy (anonymous):

\[\sqrt3+i=\sqrt{(\sqrt3)^2+1^2}\bigg(\cos\frac{\pi}{6}+i\sin\frac{\pi}{6}\bigg)=2\bigg(\cos\frac{\pi}{6}+i\sin\frac{\pi}{6}\bigg)\] and squaring gives \[(\sqrt3+i)^2=4\bigg(\cos\frac{2\pi}{6}+i\sin\frac{2\pi}{6}\bigg)\] by DeMoivre's theorem. Then \(\dfrac{2\pi}{6}=\dfrac{\pi}{3}\).

OpenStudy (anonymous):

In the complex plane: 

OpenStudy (anonymous):

Thanks a lot! I had the argument wrong...

OpenStudy (anonymous):

You're welcome